OpenStudy (anonymous):

Marlene's taxable income is $95,000. Use this tax schedule to calculate the total amount she owes in taxes. Screen Shot 2014-09-12 at 12.16.48 PM A 12,88.75 B 17,900.75 C 20,710.75 @SarahEZZMcK

OpenStudy (anonymous):

@Luigi0210

OpenStudy (texaschic101):

ok...her taxable income is 95,000 look at your chart...95,000 is over 77,100, but under 160,850. so the tax is 15,698.75 + 28% of the amount over 77,100 we need to find the amount over 77,100... 95,000 - 77,100 = 17,900 now we need to take 28% of that.....28% of 17,900 = 0.28(17,900) = 5012 now we add our 15,698.75 + 5012 = 20710.75 (C)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

dang your amazing texaschic

OpenStudy (texaschic101):

lol...I don't know about that....but thank you :)

OpenStudy (anonymous):

you are and your welcome:) imma friend you

OpenStudy (texaschic101):

sweet...friends are good

OpenStudy (anonymous):

lol yea

OpenStudy (texaschic101):

so your a country boy...lol....my neighbors are cows...hows that for country

OpenStudy (anonymous):

lol yea im a country boy well thats good but mine are horses lol

OpenStudy (anonymous):

Which of these items is a tax deduction? A cash awards B interests from bonds C interests on a home mortgage

OpenStudy (texaschic101):

interest on home mortgages are tax deductible

OpenStudy (anonymous):

last question then i will get out of your hair lol Which of these types of income is taxed? A federal tax refunds B stock dividends C interests from municipal bonds

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

I am guessing at this one....I think it is C but not 100% sure. It could be A...it is not B

OpenStudy (anonymous):

ok thanks i think i can maybe figure it out but thanks for you help for everything your awesome

OpenStudy (texaschic101):

I am going to go with A instead

OpenStudy (texaschic101):

good luck with all this tax stuff....

OpenStudy (anonymous):

ok lol that is wrong but here i need this one questions to pass Use this tax table to find how much tax you need to pay on a taxable income of $25,000. A 2,576.25 B 3,358.75 C 5,934.25

OpenStudy (texaschic101):

ok...taxable income of 25,000.... 782.50 + 15% over the amount over 7,825 25,000 - 7825 = 17175 15% of 17175 = 0.15(17175) = 2576.25 2576.25 + 782.50 = 3358.75 (B)

OpenStudy (anonymous):

Oh my gosh thank you i love you so much right now lol but not in the love love way you know but thank you so much your awesome

OpenStudy (texaschic101):

lol....glad to help :)